OpenStudy (anonymous):

Need help Determine the solution to the system of equations given below. y=x^2-5x+15 y=-x+10 A. (–2, 9) B. (2, 9) C. (2, 0) and (9, 0) D. no solution

OpenStudy (anonymous):

The solution to the system is the intersection points of the graphs of those functions. First, you have to find the x-coordinate of the intersection point. Substitute the equation 1 in equation 2 using the y-variable, so it will be like this: \(\sf x^2-5x+15 = -x+10\) Now simplify that equation and tell me what you will get for the x value.

OpenStudy (anonymous):

i got 2

OpenStudy (anonymous):

-x+10=x^2-5x+15 X^2-4x+5=0 Delta=b^2-4ac 16-4*15=-44 Delta<0 no solution Answer is D